A bag contains 6 Pink balls, 5 yellow balls and 4 red balls. Three are selected at random without replacement.
Find the probability that
(a) they are all pink
(b) two are blue and one is yellow
(c) there is one of each color .
(d) Brief your answers accordant to theories
1
Expert's answer
2020-07-27T17:54:59-0400

n=6+5+4=15 - amount number of balls

selected balls are pink:

P=(6/15)*((6-1)/(15-1))*((6-2)/(15-2))=0.044

two of selected balls are blue and one is yellow

P=0,because we have no blue balls

one of each color:

P=(C6,1*C5,1*C4,1)/C15,3=(6*5*4)/455=0.2637